THE X̃ ← Ã SEP SPECTRA OF JET-COOLED THIOPHOSGENE

\underline{PRAVEEN D. CHOWDARY}, BRENT STRICKLER, SANDRA LEE and MARTIN GRUEBELE, Department of Chemistry, University of Illinois, Urbana, IL 61801.

The X̃ ← Ã stimulated emission pumping of SCCl_{2} made over 100 new vibrational transitions accessible from 7000 – 8000 cm^{−1} above the X̃ state zero point energy.^{a} We report anharmonic constants that fit all the 482 currently assigned vibrational energy levels. A spectral autocorrelation analysis shows that the onset of free energy flow has not been reached at 8000 cm^{−1}, in accordance with higher energy B̃ state SEP data and a theoretical analysis. The available experimental data and computational results are both compatible with less fragmentation among the observable interior states of SCCl_{2}.

aP. D. Chowdary, B. Strickler, S. Lee and M. Gruebele, Chem. Phys. Lett. 434, 182 (2007).